be  was, were  been
beat  beat  beaten
become  became  become
begin  began  begun
bend  bent  bent
bite  bit  bitten
blow  blew  blown
break  broke  broken
bring  brought  brought
build  built  built
buy  bought  bought
catch  caught  caught
choose  chose  chosen
come  came  come
cost  cost  cost
cut  cut  cut
do  did  done
draw  drew  drawn
drink  drank  drunk
drive  drove  driven
eat  ate  eaten
fall  fell  fallen
feed  fed  fed
feel  felt  felt
fight  fought  fought
find  found  found
fly  flew  flown
forget  forgot  forgotten
forgive  forgave  forgiven
freeze  froze  frozen
get  got  got
give  gave  given
go  went  gone
grow  grew  grown
have  had  had
hear  heard  heard
hide  hid  hidden
hit  hit  hit
hold  held  held
hurt  hurt  hurt
keep  kept  kept
know  knew  known

lend  lent  lent
leave  left  left
let  let  let
light  lit  lit
lose  lost  lost
make  made  made
mean  meant  meant
meet  met  met
pay  paid  paid
put  put  put
ride  rode  ridden
read  read  read
ring  rang  rung
run  ran  run
say  said  said
see  saw  seen
sell  sold  sold
send  sent  sent
set  set  set
shake  shook  shaken
shine  shone  shone
shoot  shot  shot
show  showed  shown
shrink  shrank  shrunk
shut  shut  shut
sing  sang  sung
sit  sat  sat
sleep  slept  slept
speak  spoke  spoken
spend  spent  spent
split  split  split
spoil  spoilt  spoilt
stand  stood  stood
steal  stole  stolen
stick  stuck  stuck
take  took  taken
teach  taught  taught
tear  tore  torn
tell  told  told
think  thought  thought
throw  threw  thrown
understand understood understood
wake  woke  woken
wear  wore  worn
win  won  won
write  wrote  written